OpenStudy (anonymous):

Transform the following polar equation into an equation in rectangular coordinates: r = 4 A. (x - 4)2 + y2 = 4 B. x = 4 C. xy = 16 D. x2 + y2 = 16

OpenStudy (jhannybean):

Is it just r = 4? Or is there more to it...

OpenStudy (anonymous):

It's just r=4

OpenStudy (jhannybean):

Alright. Let me try working this out myself first. One second. :)

OpenStudy (jhannybean):

Alright, you understand that the equation of a circle in polar coordinates is \(x^2 +y^2 = r^2\) right?

OpenStudy (anonymous):

Right

OpenStudy (jhannybean):

Now we're already given \(r=4\), and to fit this form for the equation of a circle in polar coordinates , what would we have to do to r?

OpenStudy (anonymous):

Square it?

OpenStudy (jhannybean):

Yes! :) that means \( r^2 = 4^2 \implies r^2 = 16\)

OpenStudy (jhannybean):

Now that we've got \(r^2 = 16\) and \(x^2 +y^2 =r^2\), so which answer choice do you think it would be? :)

OpenStudy (anonymous):

Got it. Thanks!

OpenStudy (jhannybean):

Awesome! :) Good job .